how can i write the equation of a line that passes throught (-2,-1) and is perpendicular to x-2y=6?
Answer by stanbon(75887)

how can i write the equation of a line that passes throught (-2,-1) and is perpendicular to x-2y=6?
-----
Find slope of the given line:
2y = x-6
y = (1/2)x - 3
----
Slope = 1/2
-------
Slope of perpendicular line must be -2
-----
Form: y = mx + b
Solve for "b":
-1 = -2*-2+b
b = -5
-----
Equation:
y = -2x-5
================
Cheers,
Stan H.
==============